Using the expression below, choose the correct answers for the new balance and amount of interest earned in the following compound interest problem.
P(1 + r)n
$1,100 at 8%, for 15 years, compounded annually.
Total Amount = $
Interest Amount = $

Answer by Theo(13342)

the formula if f = p * (1 + r) ^ n
f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.

you have p = 1100 and r = .08 and n = 15.
the formula becomes f = 1100 * 1.08 ^ 15 = 3489.386026.
that's the total amount.
the total interest is that minus 1100 = 2389.386026.